Out of 15,000 voters in a constituency, 60% voted. Find the percentage of voters who did not vote. Can you now find how many actually did not vote?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question, it is given that
Total number of voters in the constituency = 15000
Percentage of people who voted in the election = 60%
Percentage of people who did not voted in the election = 100–60
= 40%
Total number of voters who did not voted in the election = 40% of 15000
= (40/100) × 15000 = 0.4 × 15000
= 6000 voters
∴6000 voters did not vote.

Thanks